# Silent Speech Sentence Recognition with Six-Axis Accelerometers using Conformer and CTC Algorithm

### v1

@@ Abstract
Index Terms:
Silent-speech sentence recognition, Six-axis accelerometer, Connectionist Temporal Classification (CTC), Conformer-based neural network.
I Introduction
@@
The technique for the devices mentioned above is the “silent speech interface” (SSI)[3, 4], which has been widely used to communicate with each other for people with vocal defects[5, 6]. The performance of SSI mainly depends on the techniques of silent speech signal acquisition and data decoding. The modalities of SSI used in signal acquisition include the modalities based on electromagnetic, mechanical, ultrasonic, visual, and mixed types[7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Sensors of EEG and surface electromyography (sEMG) are critical components of the electromagnetic-based SSI[17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. They are often invasive and unsuitable for improving the quality of life for the patients. Mechanical SSI uses strain and infrared sensors, while ultrasonic and visual-based SSI focuses on lip-reading from mouth shapes[27, 28, 29, 30, 31, 32, 33, 34]. Non-invasive modalities are generally complicated, expensive, impractical, and lack portability. Mixed modalities combine these techniques but often suffer from instability and unreliable signal collection due to complex structures[11, 29]. Therefore, accelerometers have been chosen for their non-invasiveness, anti-noise, portability, and high sensitivity[35, 36, 37, 38, 39].
A 95% accuracy has been achieved in word classification using three-axis accelerometers and a deep neural network[8, 40]. However, the practicality of these three-axis accelerometers is constrained by their large size and the relatively low dimensionality of the data they produce. The six-axis accelerometers provide additional gyroscope data alongside x-, y-, and z-axis data, significantly improving the portability and feature extraction.
@@
Most speech recognition approaches only consider word classification[41, 42], and not many consider continuous sentence recognition. Sentence recognition remains challenging due to elision (loss of syllables) and linking (word boundaries blending). These phenomena are especially common in English compared to Chinese or Korean [43]. The phonetic systems of Chinese and Korean feature a one-to-one correspondence between characters and phonemes, which simplifies word segmentation compared to English[43, 44]. None of the prior studies has effectively addressed the challenges of silent speech sentence recognition in dynamic and accurate segmentation and recognition (especially in English). While some researchers reported high accuracy in sentence recognition, their approaches treated entire sentences as single entities, like the word-level classification. The nuanced segmentation and dynamic contextual understanding required for accurate sentence-level recognition were overlooked[22, 45].
@@
On the other hand, various approaches based on machine learning[46, 47, 48] have been introduced to solve the segment matching problems[49]. The Conformer[50], a Transformer[51] combined with convolutional layers and Connectionist Temporal Classification (CTC) algorithms[52] has gained attention for its exceptional performance in end-to-end automatic speech recognition (ASR) systems[53]. Transformers leverage self-attention mechanisms for context understanding and the end-to-end learning, while CTC has advantages in processing sequential data without requiring strict alignment. Although CTC treats tokens as independent and lacks contextual awareness, integration with Transformers may compensate for this limitation.
@@
In this study, a new method for SSI is presented to capture the facial motion features. The proposed method employs the six-axis accelerometers together with the Conformer and CTC algorithms to decode the captured signals and to recognize sentences solely requesting the constituent words in the database. It can segment and learn dynamically from context from context to obtain high accuracy in segmenting and recognizing complex combinations of words and sentences. To the best of our knowledge, no prior study has reported the SSI method to segment and recognize sentences dynamically using a word database.
II Method
II-A Participants
@@
One laryngect and three students from Tsinghua University participated in the test to demonstrate the performance of the proposed technique. All the participants were informed of the details of our experiment before the experiment began.
Figure 1: Location of the accelerometers used to collect facial motion signals: (a) Location of sensors; (b) A patient wearing the sensor system; (c) Flowchart of the speech recognition system.
II-B Data Recording
II-B1 Microelectromechanical System
@@
The test was conducted using a microelectromechanical system, which is composed of six accelerometers and a chip ESP32 (Espressif). The communication protocol between the host and the devices is SPI, which can handle up to six devices using two buses at the same time. The accelerometer is the six-axis MPU6500 (Invensense). It records three-axis acceleration data and three-axis gyro data at the sampling frequency of 50 Hz.
@@
Fig. 1 shows the means of employing the accelerometer to record the facial motion data. Fig. 1(a) depicts the location of sensors. Fig. 1(b) shows the designed flexible printed circuit board with six accelerometers at the top of the twenty-centimeter-long “tentacles”. This design is helpful for the portability of our device and brings comfort to patients such as laryngects. Fig. 1(c) shows the flowchart of the system operation. The sensors transfer motion data to the mainboard and then to the upper monitor, at which a neural network decodes the signals and converts them into outputs.
II-B2 Data Paradigm
The accelerometers are stuck to the volunteer’s face and throat. Fig. 1(a) depicts the location of accelerometers[54, 55]. For example, channel 1 is positioned at the jaw, channel 2 is placed under the jaw, channel 3 at the throat, channel 4 and channel 5 at the upper and lower lip, respectively, while channel 6 is at the cheek.
@@ II-B2 Data Paradigm
Words and sentences in both English and Chinese were collected. Each word was recorded 100 times repeatedly, and each sample contained 80 sampling points. The size of one sample is 80 (window size) $\times$ 6 (channels) $\times$ 6 (axes). As for a sentence or phrase, the size is 180 (window size) $\times$ 6 (channels) $\times$ 6 (axes). The sentences are repeated 30 times in the experiment.
@@
In total, sixteen English words, two English sentences, and eight Chinese phrases were recorded. These words and sentences are widely used in our daily lives, especially for laryngects. To test the generalization ability of the proposed technique, ”what”, ”want”, ”wait”, ”wonder”, and ”water” are selected deliberately as easily confused words. Since test sentences contain words in the dataset, sentence recognition can be done using the existing data. All the tested words and sentences as well as their meanings are listed in Table 1.
@@
Data augmentation, such as word concatenation and random Gaussian noise, is required in that there is a strong need for sentence data to train the CTC model. Words are randomly concatenated into “sentences”, although they might be linguistically meaningless, and used for training to improve the segmentation and recognition ability of the model. The Gaussian noise with a standard deviation of one-third of the original data is added to expand the dataset size tenfold.
TABLE I: List of words, phrases, and sentences
Words Words Sentences Chinese Phrases
Afternoon What Drink water Tengtong (Hurt)
Beautiful Wait Hello, please wait Fanshen (Turn over)
Breakfast Want Xiachuang (Getting out of bed)
Drink Water Henkaixin (Very happy)
Hello Welcome Xiexieni (Thank you)
Please Wonder Woyaoheshui (I want to drink water)
Sorry Wonderful Woxiangchifan (I want to have a meal)
Thanks Tianqizhenhao (What nice weather)
II-C Data Preprocessing
@@
A slight movement by a human might be recorded in the data due to the high sensitivity of the accelerometers. Compared to facial motion features, the movements of the head and body are often in the low frequency range. To eliminate the biases and low-frequency noise, a high-pass filter is introduced. The filter is a fourth-order Butterworth filter with the cutoff frequency $f_{c}=2Hz$. For input $X$, the high-pass output $Y$ is:
@@
II-D Machine Learning Model
Sentence recognition in silent speech interfaces (SSI) is challenging due to the complexities such as elision and linking, especially in English. The complexities often make the traditional word-by-word approach ineffective. This study addresses this challenge with a transformer-based model, conformer, integrated with local multi-head self-attention. The conformer combines the ability of CNNs to capture local features with the contextual understanding of the transformer. This design enhances adaptability to varying sentence lengths and speaking rates, especially suitable for silent speech recognition.
The Connectionist Temporal Classification (CTC) algorithm addresses the challenge of aligning accelerometer data with spoken words despite variations in speech rate. By introducing a “blank” label, CTC enables the model to ignore irrelevant input frames and map multiple frames to a single output. Its loss function, based on the negative log-likelihood of the most probable output sequence, ensures robust recognition across diverse speakers. It is expressed as:
@@ II-D Machine Learning Model
II-D1 Network Structure
An end-to-end neural network based on CTC loss and Conformer with a multi-head local self-attention mechanism is used to convert accelerometer data into word sequences. Fig. 2(a) shows the complete model structure. The model mainly consists of three parts: A convolution head that processes the accelerometer data by extracting local patterns and down-sampling the data; local self-attention blocks that aggregate the context information; and the CTC algorithm that decodes the features extracted from the previous network.
Conformer is a stacked architecture combining convolutional layers and Transformer blocks. Input features pass through three convolutional blocks, each consisting of a convolution layer, batch normalization, ReLU activation, and dropout. The first block uses 30 input channels and outputs 128 hidden dimensions. A residual connection is applied in the second block, while the third block down-samples data to enhance computational efficiency.
Local self-attention was applied which enforces the attention matrix to be a band matrix. This makes each token only obtain information from a fixed-size surrounding window (which is about 0.4s in this work). To maintain the model’s ability to gain access to the global context, a global token at the beginning of each sequence is appended to aggregate overall information.
The CTC algorithm is introduced for its excellent performance in sequence-to-sequence tasks. CTC works by segmenting the words into thinner slices (smaller window size) and recognizing them as well as considering the entire sequence of inputs and outputs, rather than requiring a one-to-one alignment between input frames and output labels.
II-D2 Training Parameters
The model was trained to minimize the sum of CTC loss and Cross-Entropy loss, using the Adams optimizer. The learning rate was set to 0.001, with a decay rate of 0.00001. Eighty percent of the dataset was used for blind training, and the remaining twenty percent was used for testing. The batch size was 32 and the number of epochs was set to 50. PyTorch is used for all data training processes of the neural network mentioned above.
III Results
III-A Accuracy of Recognition
The proposed sentence recognition method trains the model to learn from sentence data and then uses the existing words to identify each part of an unknown sentence.
@@ III-A Accuracy of Recognition
The model shows an excellent performance in sentence segmentation and recognition based on the good quality of the accelerometer data. The Conformer performs well in extracting the features and learning the contextual information. The CTC algorithm segments the sentence and decodes the result. Accuracy is calculated as the ratio of correctly recognized words to the total words in a sentence or dataset. Fig. 3(c) shows the accuracy and the loss curve during the training process. It can be seen from Fig. 3(c) that the total accuracy in the whole dataset is 97.17%, which is higher than previous studies [8], which showed an accuracy of 94.65$\pm$2.54%. As shown in Fig. 3(c), the average classification accuracy is 97.24% for words and 97.16% for the sentences.
III-B Performance of Easily Confused Words
@@
The classification accuracy on five easily confused words (what, want, wait, wonder, and water) is analyzed to evaluate the discriminative ability of the proposed method in distinguishing phonetically similar words. These words were chosen due to their high potential for misclassification, given their resemblance in articulation and motion patterns.
@@
Fig. 3(d) is the confusion matrix of the five easily confused words. It offers a visual representation of how the model differentiates the easily confused words. It can be seen from Fig. 3(d) that the model performs well in classifying these ambiguous words, achieving a high level of precision in distinguishing their subtle differences in facial motion and pronunciation. The mean accuracy exceeds 99%, demonstrating the robustness of the system. Results also showed that only roughly 0.2% of the testing samples were incorrectly recognized. This implies that misclassification occurs at a very low rate. This impressive performance highlights the effectiveness of the accelerometer-based system combined with the sequence-to-sequence algorithm. The results suggest that the proposed approach not only captures gross facial movements and mouth shape variations but also extracts fine-grained motion features from the accelerometer data. Such capability is crucial for improving the accuracy of silent speech recognition and ensuring that highly similar words can be correctly classified with minimal errors.
III-C Performance of Sentence Recognition
@@
In addition to word classification, the proposed model has demonstrated excellent performance in sentence segmentation and recognition, further validating its capability in processing continuous speech data. The average sentence recognition accuracy reaches 97.16%, indicating a high level of reliability in translating facial motion patterns into corresponding textual representations. This superior performance is largely attributed to the Conformer model’s ability to effectively capture both local dependencies and global contextual information. By integrating self-attention mechanisms with convolutional layers, the model ensures that even subtle variations in facial motion are accurately mapped to words and sentences, thereby improving the overall recognition precision.
@@
Fig. 4(a) illustrates the similarity between a portion of sentence data and the corresponding individual words in the original time domain. This visualization confirms that the model successfully maintains structural integrity and consistency when processing longer speech sequences. Additionally, Fig. 4(c) presents the relationship between the average recognition accuracy and sentence length. The results show a relatively stable trend, with accuracy consistently remaining above 95% across different sentence lengths. This demonstrates that the use of CTC decoding effectively mitigates the impact of sentence length on recognition performance. Unlike traditional methods that may struggle with longer sequences, the CTC-based approach enables the model to maintain high accuracy regardless of sentence complexity.
III-D Performance with respect to Different Individuals
@@
The model’s performance across different participants is illustrated in Fig. 4(b). The red bars in the figure represent the accuracy achieved by each participant under standard conditions, while the green bars depict “blind” testing results. In the “blind” scenario, the model was trained using data from other participants and tested on a new individual without prior exposure to their specific motion patterns. This evaluation was conducted to assess the generalization capability of the model when applied to unseen subjects.
The results indicate that, on average, participants achieved over 95% accuracy when the model was trained using their own data, demonstrating its effectiveness in adapting to individual speech patterns. However, the “blind” accuracy varied among individuals, reflecting differences in speaking habits, facial muscle movements, and pronunciation styles. While some participants achieved relatively high recognition rates even in the blind testing scenario, others exhibited a noticeable drop in accuracy due to unique articulation characteristics.
@@
To address this variability and enhance the model’s generalization ability, a “few-shot” learning approach was applied. By incorporating a small number of participant-specific samples into the training dataset, the recognition accuracy was significantly improved. The blue bars in Fig. 4(b) indicate that with only a limited amount of personalized data, the model’s accuracy increased to over 90%, effectively reducing the performance gap observed in blind testing. This result demonstrates the model’s capacity to adapt to individual differences and highlights the importance of incorporating minimal personalized training data to achieve more reliable recognition results across diverse users.
III-E Performance with respect to Different Channels

### v2

@@ Abstract
{IEEEkeywords} Silent speech sentence recognition, six-axis accelerometer, Connectionist Temporal Classification (CTC), Conformer-based neural network.
1 Introduction
[…]
@@ 1 Introduction
The technique for the devices mentioned above is the “silent speech interface” (SSI) [3, 4], which has been widely used to communicate with each other for people with vocal defects [5, 6]. The performance of SSI mainly depends on the techniques of silent speech signal acquisition and data decoding. The modalities of SSI used in signal acquisition include the modalities based on electromagnetic, mechanical, ultrasonic, visual, and mixed types [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Sensors of EEG and surface electromyography (sEMG) are critical components of the electromagnetic-based SSI [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. They are usually bulky systems and are not portable , which makes them unsuitable for improving the quality of life for the patients. Mechanical SSI mainly uses strain sensors, while ultrasonic and visual-based SSI focuses on lip-reading from mouth shapes [27, 28, 29, 30, 31, 32, 33, 34]. Non-invasive modalities are generally complicated, expensive, and some are even impractical, and lack portability [2]. Mixed modalities combine these techniques but often suffer from instability due to complex structures [11, 29]. Therefore, accelerometers have been chosen for their non-invasiveness, anti-noise, portability, and high sensitivity [35, 36, 37, 38, 39].
@@
Most speech recognition approaches only consider word classification [8, 40, 41, 42], and not many consider continuous sentence recognition. Here, sentence recognition refers to the ability to correctly translate an unfamiliar long sentence into a sequence of words based on an existing word database. Sentence recognition remains challenging due to elision (loss of syllables) and linking (word boundaries blending). These phenomena are especially common in English compared to Chinese or Korean [43]. The phonetic systems of Chinese and Korean feature a one-to-one correspondence between characters and phonemes, which simplifies the word segmentation compared to English [43, 44]. None of the prior studies has effectively addressed the challenges of silent speech sentence recognition in dynamic and accurate segmentation and recognition (especially in English). While some researchers reported high accuracy in sentence recognition, their approaches treated entire sentences as single entities, just like word-level classification. The nuanced segmentation and dynamic contextual understanding required for accurate sentence-level recognition were overlooked [22, 45].
@@
On the other hand, various approaches based on machine learning [46, 47, 48] have been introduced to solve segment-matching problems [49]. The Conformer [50], a Transformer [51] combined with convolutional layers and Connectionist Temporal Classification (CTC) algorithms [52], has gained attention for its exceptional performance in end-to-end automatic speech recognition (ASR) systems [53]. Transformers leverage self-attention mechanisms for context understanding and the end-to-end learning, while CTC has advantages in processing sequential data without requiring strict alignment. Although CTC treats tokens as independent and lacks contextual awareness, integration with Transformers may compensate for this limitation.
@@
In this study, a new method for SSI is presented to capture the facial motion features. The proposed method employs the six-axis accelerometers together with the Conformer and CTC algorithm to decode the captured signals and to recognize sentences solely requesting the constituent words in the database. Instead of treating the whole sentence as an entire pattern, the system can learn dynamically from context from context to obtain high accuracy in segmenting and recognizing complex combinations of words and sentences. To the best of our knowledge, no prior study has reported the SSI method to segment and recognize sentences dynamically.
2 Method
2.1 Participant
@@
One laryngect and three undergraduates participated in the test to evaluate the performance of the proposed method. All the participants were informed of the details of the experiment before the experiment began.
2.2 Data Recording
2.2.1 Microelectromechanical System
@@
The test was conducted using a microelectromechanical system, which is composed of six six-axis accelerometers and a chip ESP32 (Espressif). The communication protocol between the host and the devices is SPI, which can handle up to six devices using two buses at the same time. The accelerometer is the six-axis MPU6500 (Invensense). It records three-axis acceleration data and three-axis gyro data at the sampling frequency of 50 Hz.
Figure 1: Location of the accelerometers used to collect facial motion signals: (a) Location of sensors; (b) A patient wearing the sensor system
@@
Fig. 1 shows the means of employing the accelerometer to record the facial motion data. Fig. 1(a) depicts the location of accelerometers[54, 55]. Fig. 1(b) shows the designed flexible printed circuit board with six accelerometers at the top of the twenty-centimeter-long “tentacles”. This design is helpful for the portability of our device and brings comfort to patients such as laryngects.
2.2.2 Data Paradigm
@@
Words and sentences in both English and Chinese were collected. Each word was recorded 100 times repeatedly, and each sample contained 80 sampling points, so the size of one sample is 80 (window size) $\times$ 6 (channels) $\times$ 6 (axes). As for a sentence or phrase, the window size is 180 and is repeated 30 times in the experiment. The system doesn’t treat sentences as single patterns, or mix words and sentences together for classification. Instead, it is learning the characteristics from the words as well as the associations and connections of the motions between those words from the sentences. Therefore, sentences containing multiple words have a larger window size.
@@
In total, sixteen English words, two English sentences, and eight Chinese phrases that are widely used in daily lives were recorded. To test the generalization ability of the proposed technique, ”what”, ”want”, ”wait”, ”wonder”, and ”water” are selected deliberately as easily confused words. All the tested words and sentences as well as their meanings are listed in Table 1.
Table 1: List of Words, Sentences, and Phrases
Words Afternoon, Thanks, Beautiful, Wait, Breakfast, Want, Drink, Water, Hello, Welcome, Please, What, Sorry, Wonder, Test, Wonderful
Sentences Drink water. Hello, please wait.
Chinese Phrases Tengtong (Hurt), Fanshen (Turn over), Xiachuang (Getting out of bed), Henkaixin (Very happy), Xiexieni (Thank you), Woyaoheshui (I want to drink water), Woxiangchifan (I want to have a meal), Tianqizhenhao (What nice weather)
@@
The whole dataset is randomly split into training, validation, and testing dataset with ratio 70%, 15%, and 15% for future training. Data augmentation, such as word concatenation and random Gaussian noise, is also required for training dataset because there is a strong need for sentence data to train the CTC model. Words are randomly concatenated into “sentences”, although they might be linguistically meaningless, and used for training to improve the segmentation and recognition ability of the model. The Gaussian noise with a standard deviation of one-third of the original data is added to expand the dataset size tenfold.
2.3 Data Preprocessing
@@
A slight movement by a human (motion artifacts) might be recorded in the data due to the high sensitivity of the accelerometers. Compared to facial motion features, the movements of the head and body are often in the low frequency range. To eliminate the biases and low-frequency noise, a high-pass filter is introduced. The filter is a fourth-order Butterworth filter with the cutoff frequency $f_{c}=2Hz$. For input $X$, the high-pass output $Y$ is:
@@
2.4 Machine Learning Model
2.4.1 Network Structure
Sentence recognition in silent speech interfaces (SSI) is challenging due to the complexities such as elision and linking, especially in English. The complexities often make the traditional word-by-word approach ineffective. This study addresses this challenge with an end-to-end neural network based on Conformer with a multi-head local self-attention mechanism and CTC loss. Fig. 2(a) shows the complete model structure. The model mainly consists of three parts: A convolution head that processes the accelerometer data by extracting local patterns and down-sampling the data; local self-attention blocks that aggregate the context information; and the CTC algorithm that decodes the features extracted from the previous network.
Conformer is a stacked architecture combining convolutional layers and Transformer blocks. Input features pass through three convolutional blocks, each consisting of a convolution layer, batch normalization, ReLU activation, and dropout. The first block uses 30 input channels and outputs 128 hidden dimensions. A residual connection is applied in the second block, while the third block down-samples data to enhance computational efficiency. Local self-attention was applied, which enforces the attention matrix to be a band matrix, as shown in Fig. 2(b). This makes each token only obtain information from a fixed-size surrounding window. To maintain the model’s ability to gain access to the global context, a global token at the beginning of each sequence is appended to aggregate overall information.
The Connectionist Temporal Classification (CTC) algorithm performs well in sequence-to-sequence tasks. It addresses the challenge by introducing a “blank” label. CTC enables the model to ignore irrelevant input frames and map multiple frames to a sequence of outputs. Its loss function is expressed as:
@@ 2.4.1 Network Structure
2.4.2 Training Parameters
The model was trained to minimize the sum of CTC loss and Cross-Entropy loss, using the Adams optimizer. The learning rate was set to 0.001, with a decay rate of 0.00001. The batch size was 32 and the number of epochs was set to 50. PyTorch is used for all data training processes of the neural network mentioned above.
3 Results
3.1 Accuracy of recognition
@@
The model shows an excellent performance in sentence segmentation and recognition. Accuracy is calculated as the ratio of correctly recognized words to the total words in a sentence or dataset. Fig. 3(c) shows the accuracy and the loss curve on the validation dataset during the training process. It can be seen from Fig. 3(c) that the accuracy is 97.17%, which is higher than previous studies [8], which showed an accuracy of 94.65±2.54%. As shown in Fig. 3(c), the average classification accuracy is 97.24% for words and 97.16% for the sentences.
3.2 Performance of Easily Confused Words
@@
The classification accuracy on five easily confused words (“what”, “want”, “wait”, “wonder”, and “water”) is analyzed to evaluate the discriminative ability of the proposed method in distinguishing phonetically similar words. These words were chosen due to their high potential for misclassification, given their resemblance in articulation and motion patterns.
@@
Fig. 3(d) is the confusion matrix of the five easily confused words. The mean accuracy exceeds 99%, demonstrating the robustness of the system. Results also showed that only roughly 0.2% of the testing samples were incorrectly recognized. This implies that misclassification occurs at a very low rate. The results suggest that the proposed approach not only captures gross facial movements and mouth shape variations but also extracts fine-grained motion features from the accelerometer data.
3.3 Performance of Sentence Recognition
@@
In addition to word classification, the proposed model has demonstrated excellent performance in sentence segmentation and recognition, further validating its capability in processing continuous speech data. The average sentence recognition accuracy reaches 97.16%, indicating a high level of reliability in translating facial motion patterns into corresponding textual representations. This superior performance is largely attributed to the Conformer model’s ability to effectively capture both local dependencies and global contextual information.
@@
Fig. 4(a) illustrates the similarity between a portion of sentence data and the corresponding individual words in the original time domain. This visualization confirms that the model successfully maintains structural integrity and consistency when processing longer speech sequences. Additionally, Fig. 4(c) presents the relationship between the average recognition accuracy and sentence length. The results show a relatively stable trend, with accuracy consistently remaining above 95% across different sentence lengths. This demonstrates that unlike traditional methods, the CTC decoding effectively mitigates the impact of sentence length on recognition performance.
3.4 Performance with respect to Different Individuals
@@
The model’s performance across different participants is illustrated in Fig. 4(b). The red bars in the figure represent the accuracy achieved by each participant under standard conditions, while the green bars depict “blind” testing results. In the “blind” scenario, the model was trained using data from other participants and tested on a new individual without prior exposure to their specific motion patterns. This evaluation was conducted to assess the generalization capability of the model when applied to unseen subjects. The results indicate that, on average, participants achieved over 95% accuracy when the model was trained using their own data, demonstrating its effectiveness in adapting to individual speech patterns. However, the “blind” accuracy varied among individuals, reflecting differences in speaking habits, facial muscle movements, and pronunciation styles.
@@
To address this variability and enhance the model’s generalization ability, a “few-shot” learning approach was applied. By incorporating a small number of participant-specific samples into the training dataset, the recognition accuracy was significantly improved. The blue bars in Fig. 4(b) indicate that with only a limited amount of personalized data, the model’s accuracy increased to over 90%, effectively reducing the performance gap observed in blind testing. This result demonstrates the model’s capacity to adapt to individual differences.
3.5 The performance with respect to Different Channels